suppose mr arsalan's basic salary is 12000.calculate the allowances (house rent,utility bills,conveyance allowanc).Also,find the provident fund and gratuity?
1
Expert's answer
2019-11-22T10:31:32-0500

Basic Salary = 12,000


Regardless, we need to discover every stipend aggregate. 


House Rent: As v acknowledge house rent is 45% of basic compensation along these lines, 


12,000 x 45% = 5400


Administration Bills: As v acknowledge Utility is 2.5% of principal remuneration along these lines, 


12,000 x 2.5% = 300

Development stipend: As v acknowledge transport is 2.5% of fundamental compensation along these lines, 


12,000 x 2.5% = 300


Directly outright settlement = house rent + Utility Bills + Conveyance stipend 


= 5400 + 300 +300 


Hard and fast ALLOWANCE = 6000


By and by finding Provident hold aggregate and tip, 


Lucky Fund: V all acknowledge Provident support is 1/eleventh of basic pay so in this circumstance, 


Blessed support = 1/11 x 12,000 


= 1091/ - yet v moreover understand that identical proportion of Provident hold furthermore contributed by the association along these lines, 1091+1091 = 2181


Fortunate finance = 2181


Tip Fund: It's in like manner 1/eleventh of fundamental remuneration along these lines, 


Tip account = 1/11x 12,000 


Tip sponsor = 1091